An inductor, a capacitor, and a resistor are all connected in series across an ac source.

If the resistance, inductance, and capacitance are all doubled, by what factor does each of the

following quantities change


1
Expert's answer
2022-03-21T12:46:53-0400

LCR series combination

Impedance

"Z=\sqrt{R^2+(X_c-X_L)}"

"X_L=X_c"

Resonance frequency

"wL=\frac{1}{wC}"

"w=\frac{1}{\sqrt{LC}}"

L=2L

C=2C

"w'=\frac{1}{\sqrt{4LC}}"

"\frac{w'}{w}=\frac{1}{2}"

"w'=\frac{w}{2}"

Inductive reactance

"X_L=wL"

"X'_L={2wL}\\L'=2L"

"\frac{X'_L}{X_L}=\frac{2w_L}{wL}"

"X'_L=2X_L"

Capacitive reactance

"X_c=\frac{1}{wC}"

C'=2C

"X'_c=\frac{1}{2wC}"

"\frac{X'_c}{X_c}=\frac{1}{2}"

"X'_C=\frac{X_c}{2}"